OpenStudy (anonymous):

how do i verify the identity cos 2 theta= 2 cos^2-i using an angel sum idenrtity

OpenStudy (anonymous):

Proof for: \(\cos2\theta=2\cos^2\theta-1\): \(\cos 2\theta=\cos^2\theta-\sin^2\theta\) ** \(=\cos^2\theta-(1-\sin^2\theta)\) \(=2\cos^2\theta-1\) **Remember: \(cos(a+b)=\cos a \cos b-\sin a \sin b\)

OpenStudy (anonymous):

epic thank you